determine the test statistic, critical values, and the p-value, 90% confidence interval estimateA survey of 700 adults from a certain region asked, If purchasing a used car made certain upgrades or features more affordab
Answer #1

pcap = ( 236+ 159)/(400 +300)
= 0.5643


zstat = (p1 -p2)/sqrt(pcap *(1-pcap) *(1/n1+1/n2))
= ( 0.59 -0.53)/sqrt(0.5643 *(1-0.5643) *(1/400 +1/300))
= 1.58


Critical value = +/- 1.645

p value = 0.1131


Do not reject H0

z value at 90% = 1.645

CI = (p1 -p2) +/- z *sqrt(p1*(1-p1)/n1+p2*(!-p2)/n2)
= (0.59 -0.53) +/- 1.645 *sqrt(0.59 *(1-0.59)/400 + 0.53 *(1-0.53)/300)

= (-0.0023 , 0.1223)